Reduce the following equations to standard form, and then identify which conicoids they represent. Further, give a rough sketch of each.
i) x^2 +y^2 +2x−y−z+3 = 0
ii) 3y^2 +3z^2 +4x+3y+z = 9
1
Expert's answer
2020-03-13T10:13:39-0400

"Given \\ equation\\ is \\ x^2+y^2+2x-y-z+3\\\\\nCompleting \\ the \\ squares \\ we \\ get \\ \nx^2+2x+1+y^2-2(\\frac{1}{2})(y)+\\frac{1}{4}-\\frac{1}{4}-z-\\frac{1}{4}+2=0\\\\\n\\Rightarrow\\ (x+1)^2+(y-\\frac{1}{2})^2+\\frac{7}{4}-z=0\\\\\n\\Rightarrow \\ z=(x+1)^2+(y-\\frac{1}{2})^2+\\frac{7}{4}\\\\\nThe\\ graph\\ of \\ the \\ given\\ equation\\ is\n\u200b\t\n\n \n \n\u200b\t\n\n\u200b"



"ii) \ 3y^2+3z^2+4x+3y+z=9\\
Completing\ the\ squares \ we\ get\\
3(y+\frac{1}{2})^2+3(z+\frac{1}{6})^2+4x=(\sqrt\frac{59}{6})^2\\
The\ graph \ of \ the\ given\ curve\ is \\"